OpenStudy (anonymous):

the measure of angle k is 16 degrees less than three times the measure of its supplement

OpenStudy (anonymous):

tht is notthe answer

OpenStudy (anonymous):

x=(3(180-x))-16 x=540-3x-16 4x=524 x=131